# Transferable Reinforcement Learning via Generalized Occupancy Models

### v2

@@ Abstract
Intelligent agents must be generalists, capable of quickly adapting to various tasks. In reinforcement learning (RL), model-based RL learns a dynamics model of the world, in principle enabling transfer to arbitrary reward functions through planning. However, autoregressive model rollouts suffer from compounding error, making model-based RL ineffective for long-horizon problems. Successor features offer an alternative by modeling a policy’s long-term state occupancy, reducing policy evaluation under new tasks to linear reward regression. Yet, policy improvement with successor features can be challenging. This work proposes a novel class of models, i.e., generalized occupancy models (GOMs), that learn a distribution of successor features from a stationary dataset, along with a policy that acts to realize different successor features. These models can quickly select the optimal action for arbitrary new tasks. By directly modeling long-term outcomes in the dataset, GOMs avoid compounding error while enabling rapid transfer across reward functions. We present a practical instantiation of GOMs using diffusion models and show their efficacy as a new class of transferable models, both theoretically and empirically across various simulated robotics problems. Videos and code: https://weirdlabuw.github.io/gom/.
@@ 1 Introduction
Figure 1: The transfer setting for generalized occupancy models. Given an unlabeled offline dataset, we learn a generalized occupancy model that models both “what can happen?" $p(\psi|s)$ and “how can we achieve a particular outcome?" $p(a|s,\psi)$. This is used for quick adaptation to new downstream tasks without test-time policy optimization.
@@
This work addresses the challenge of developing RL agents that can broadly generalize to any task in an environment specified by a reward function. To achieve this type of generalization, we consider the paradigm of pretraining on an offline dataset of transitions and inferring optimal policies for downstream tasks from observing task-specific rewards. Since the target task is not revealed during pretraining, the model must encode information about the environment dynamics without committing to a particular policy or reward. Moreover, once the task reward is observed, the model must provide a way to evaluate and improve the policy since different tasks require different optimal policies.
@@
A natural approach to this problem is model-based reinforcement learning [60, 20, 61], which learns an approximate dynamics model of the environment. Given a downstream reward function, task-optimal behavior can be obtained by “planning” via model rollouts [49, 57, 38, 44]. Typically, model rollouts are generated in an autoregressive fashion, conditioning each step on generation from the previous steps. In practice, however, autoregressive generation notoriously suffers from compounding error [30, 1, 25], which arises when small, one-step approximation errors accumulate over long horizons. This leads to rollout trajectories that diverge from real trajectories, limiting many model-based RL methods to short-horizon, low-dimensional problems.
@@
A alternative class of algorithms based on successor features (SFs) has emerged as a potential approach to transferable decision-making [3, 2]. Successor features represent the discounted sum of features for a given policy. Assuming a linear correspondence between features and rewards, policy evaluation under new rewards reduces to a simple linear regression problem. Notably, by directly predicting long-term outcomes, SFs avoid autoregressive rollouts and hence compounding error. However, the notion of successor features is deeply tied to the choice of a particular policy. This policy dependence poses a challenge when recovering the optimal policy for downstream tasks. Current approaches to circumvent policy dependence either maintain a set of policies and select the best one during inference [3] or randomly sample reward vectors and make conditional policy improvements [5, 53, 54]. Nevertheless, a turnkey solution to transfer remains a desirable goal.
@@
In this work, we propose a new class of models—generalized occupancy models (GOMs)—that are rapidly transferable across reward functions while avoiding compounding errors. Rather than modeling the successor features under a particular policy, GOMs model the entire distribution of successor features under the behavior policy, effectively encoding all possible outcomes that appear in the dataset starting from a designated state. Crucially, by representing outcomes as successor features, we enjoy the benefit of zero-shot outcome evaluation after solving a linear reward regression problem, while avoiding compounding error. To enable transfer to downstream tasks, we jointly learn a readout policy that generates an action to accomplish a particular long-horizon outcome (represented as a successor feature) from a state. At test time, we plan by selecting the best in-distribution outcome from the current state for a particular reward and query the readout policy for an optimal action to realize it. GOMs are a new class of world models as they essentially capture the dynamics of the world and can be used to plan for arbitrary rewards in new tasks.
@@
Since multiple outcomes can follow from a particular state, and multiple actions can be taken to achieve a particular outcome, both the outcome distribution and the policy require expressive model classes to represent. We provide a practical instantiation of GOMs using diffusion models [22, 46] and show that under this parameterization, planning can be cast as a simple variant of guided diffusion sampling [12]. We validate the transferability of GOMs across a suite of long-horizion simulated robotics domains and further provide analysis showing that GOMs provably converge to “best-in-data” policies. With GOMs, we hope to introduce a new way for the research community to envision transfer in reinforcement learning beyond model-based RL.
@@ 2 Related Work
Model-Based RL To enable transfer across rewards, model-based RL learns one-step (or multi-step) dynamics models via supervised learning and use them for planning  [10, 37, 38, 57] or policy optimization  [11, 49, 25, 63, 19]. These methods typically suffer from compounding error, where autoregressive model rollouts lead to large prediction errors over time  [1, 30]. Despite improvements to model architectures  [19, 24, 1, 31, 64] and learning objectives [26], modeling over long horizons without compounding error remains an open problem. GOMs instead directly model cumulative long-term outcomes in an environment, avoiding autoregressive generation while remaining transferable.
@@
Successor Features Successor features achieve generalization across rewards by modeling the accumulation of features (as opposed to rewards in model-free RL)  [3, 2]. With the assumption that rewards are linear in features, policy evaluation under new rewards reduces to a linear regression problem. A key limitation of successor features is their inherent dependence on a single policy, as they are defined as the accumulated features when acting according to a particular policy. This makes extracting optimal policies for new tasks challenging.
@@
To circumvent this policy dependence, generalized policy improvement [3, 2] maintains a discrete set of policies and selects the highest valued one to execute at test time, limiting the space of available policies for new tasks. Universal SF [6] and forward-backward representations [53, 54] randomly sample reward weights $z$ and jointly learn successor features and policies conditioned on $z$. The challenge lies in achieving coverage over the space of all possible policies through sampling of $z$, resulting in potential distribution shifts for new problems. RaMP [8] learns a successor feature predictor conditioned on an initial state and a sequence of actions. Transfer requires planning by sampling actions sequences, which becomes quickly intractable over horizon. In contrast, GOMs avoid conditioning on any explicit policy representation by modeling the distribution of all possible outcomes represented in a dataset, and then selecting actions corresponding to the most desirable long term outcome.
@@
Distributional Successor Measure (DSM) [58] is a concurrent work that learns a distribution over successor representations using tools from distributional RL [4]. Importantly, DSM models the distributional successor measure of a particular policy, where the stochasticity stems purely from the policy and the dynamics. This makes it suitable for robust policy evaluation but not for transferring to arbitrary downstream tasks. In contrast, GOMs model the distribution of successor feature outcomes in the dataset (i.e., the behavior policy). Here the distribution stems from range of meaningfully distinct long-term outcomes. This type of modeling allows GOMs to extract optimal behvavior for arbitrary downstream tasks, while DSMs suffer from the same policy dependence that standard successor feature-based methods do.
@@ 3 Preliminaries
Successor features [2] generalize the notion of a value function from task-specific rewards to task-agnostic features. Given a state feature function $\phi:S\rightarrow\mathbb{R}^{d},$ the successor feature of a policy is defined as $\psi^{\pi}(s)=\mathbb{E}_{\pi,\mathcal{T}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}\phi(s_{i})\mid s_{1}=s\right]$. Suppose rewards can be linearly expressed by the features, i.e. there exists $w\in\mathbb{R}^{n}$ such that $R(s)=w^{\top}\phi(s),$ then the value function for the particular reward can be linearly expressed by the successor feature $V^{\pi}(s)=w^{\top}\psi^{\pi}(s).$ Hence, given the successor feature $\psi^{\pi}$ of a policy $\pi$, we can immediately compute its value under any reward once the reward weights $w$ are known. Analogous to value functions, successor features also admit a recursive Bellman identity $\psi^{\pi}(s):=\phi(s)+\gamma\mathbb{E}_{\pi}\left[\psi^{\pi}(s^{\prime})\right],$ allowing them to be estimated using dynamic programming [3]. In this paper, with a slight abuse of terminology, we also refer to the discounted sum of features along a trajectory as a successor feature. In this sense, a successor feature represents an outcome that is feasible under the dynamics and can be achieved by some policy.
Diffusion Models GOMs rely on expressive generative models to represent the distribution of successor features. Diffusion models [22, 46] are a class of generative models where data generation is formulated as an iterative denoising process. Specifically, DDPM [22] consists of a forward process that iteratively adds Gaussian noise to the data, and a corresponding reverse process that iteratively denoises a unit Gaussian to generate samples from the data distribution. The reverse process leverages a neural network estimating the score function of each noised distribution, trained with a denoising score matching objective [48]. In addition, one can sample from the conditional distribution $p(x|y)$ by adding a guidance $\nabla_{x}\log p(y|x)$ to the score function in each sampling step [12]. As we show in Sec. 4.3, guided diffusion enables quick sampling of optimal outcomes from GOMs.
@@
4 Generalized Occupancy Models
We introduce the framework of GOMs as a scalable approach to the transfer problem described in Section 3, with the goal of learning from an unlabeled dataset to quickly adapt to any downstream task specified by a reward function. We start by relating the technical details behind learning GOMs in Section 4.1, followed by explaining how GOMs can be used for efficient multi-task transfer in Section 4.2. Finally, we describe a practical instantiation of GOMs in Section 4.3.
4.1 Learning Generalized Occupancy Models
To transfer and obtain optimal policies across different reward functions, generalist decision-making agents must model the future in a way that permits the evaluation of new rewards and new policies. To this end, GOMs adopt a technique based on off-policy dynamic programming to directly model the distribution of cumulative future outcomes, without committing to a particular reward function $r(\cdot)$ or policy $\pi$. Fig. 2 illustrates the two components in GOMs, and we describe each below.
(1) Outcome model: for a particular a state feature function $\phi(s)$, GOMs model the distribution of successor features $p(\psi|s)$ over all paths that have coverage in the dataset. In deterministic MDPs, each successor feature $\psi$ (discounted sum of features $\psi=\sum_{t}\gamma^{t-1}\phi(s_{t})$) can be regarded as an “outcome". When the state features are chosen such that reward for the desired downstream task is a linear function of features, i.e., there exists $w\in\mathbb{R}^{n}$ such that $r(s)=w^{\top}\psi(s)$  [41, 42, 65, 55, 8, 3], the value of each outcome can be evaluated as $w^{\top}\psi$. That is, knowing $w$ effectively transforms the distribution of outcomes $p(\psi|s)$ into a distribution of task-specific values (sum of rewards) $p(R|s)$. Notably, since $w$ can be estimated by regressing rewards from features, distributional evaluation on a new reward function boils down to a simple linear regression.
As in off-policy RL, the outcome distribution $p(\psi|s)$ in GOMs can be learned via an approximate dynamic programming update, which is similar to a distributional Bellman update [4]:
@@ 4.1 Learning Generalized Occupancy Models
(2) Readout policy: Modeling the distribution of future outcomes in an environment is useful only when it can be realized in terms of actions that accomplish particular outcomes. To do so, GOMs pair the outcome model with a readout policy $\pi(a|s,\psi)$ that actualizes a desired long-term outcome $\psi$ into the action $a$ to be taken at state $s$. Along with the outcome model $p_{\theta}(\psi|s)$, the readout policy $\pi_{\rho}(a|s,\psi)$ can be optimized via maximum-likelihood estimation:
@@
Figure 2: Generalized Occupancy Model for a simple environment. Given a state feature function $\phi$, GOMs learn a distribution of all possible long-term outcomes (successor features $\psi$) in the dataset $p(\psi|s)$, along with a readout policy $\pi(a|s,\psi)$ that takes an action $a$ to realise $\psi$ starting at state $s$.
@@
4.2 Transfer with Generalized Occupancy Models
To synthesize optimal policies for novel downstream reward functions using GOMs, two sub-problems must be solved: (1) inferring the suitable linear reward weights $w_{r}$ for a particular reward function from a set of $(s,r)$ tuples and (2) using the inferred $w_{r}$ to select an optimal action $a^{*}$ at a state $s$. We discuss each below.
Inferring task-specific weights with linear regression. As noted, for any reward function $r(s)$, once the linear reward weights $w_{r}$ are known (i.e., $r(s)=w_{r}^{T}\phi(s)$), the distribution of returns in the dataset $p(R|s)$ is known through linearity. However, in most cases, rewards are not provided in functional form, making $w_{r}$ unknown a priori. Instead, given a dataset of $\mathcal{D}=\{(s,r)\}$ tuples, $w_{r}$ can be obtained by solving a simple linear regression problem $\min\frac{1}{|\mathcal{D}|}\sum_{(s,r)\in\mathcal{D}}\|w_{r}^{T}\psi(s)-r\|_{2}^{2}$. We emphasize that this is all the optimization required for transferring GOMs to new tasks.
Generating task-specific policies via distributional evaluation. Given the inferred $w_{r}$ and the corresponding future return distribution $p(R|s)$ obtained through linear scaling of $p(\psi|s)$, the optimal action can be obtained by finding the $\psi$ corresponding to the highest possible future return that has sufficient data-support:

### v3

# Distributional Successor Features Enable Zero-Shot Policy Optimization

@@ Abstract
Intelligent agents must be generalists, capable of quickly adapting to various tasks. In reinforcement learning (RL), model-based RL learns a dynamics model of the world, in principle enabling transfer to arbitrary reward functions through planning. However, autoregressive model rollouts suffer from compounding error, making model-based RL ineffective for long-horizon problems. Successor features offer an alternative by modeling a policy’s long-term state occupancy, reducing policy evaluation under new rewards to linear regression. Yet, policy optimization with successor features can be challenging. This work proposes a novel class of models, i.e., Distributional Successor Features for Zero-Shot Policy Optimization (DiSPOs), that learn a distribution of successor features of a stationary dataset’s behavior policy, along with a policy that acts to realize different successor features within the dataset. By directly modeling long-term outcomes in the dataset, DiSPOs avoid compounding error while enabling a simple scheme for zero-shot policy optimization across reward functions. We present a practical instantiation of DiSPOs using diffusion models and show their efficacy as a new class of transferable models, both theoretically and empirically across various simulated robotics problems. Videos and code are available at https://weirdlabuw.github.io/dispo/.
@@ 1 Introduction
Figure 1: The transfer setting for DiSPOs. Given an unlabeled offline dataset, DiSPOs model both “what can happen?" $p(\psi|s)$ and “how can we achieve a particular outcome?" $p(a|s,\psi)$. This is used for quick adaptation to new downstream tasks without test-time policy optimization.
@@
This work addresses the challenge of developing RL agents that can broadly generalize to any task in an environment specified by a reward function. To achieve this type of generalization, we consider the paradigm of pretraining on an offline dataset of transitions and inferring optimal policies for downstream tasks from observing task-specific rewards. Since the target task is not revealed during pretraining, the model must encode information about the environment dynamics without committing to a particular policy or reward. Moreover, once the task reward is observed, the model must provide a way to quickly evaluate and improve the policy since different tasks require different optimal policies.
@@
A natural approach to this problem is model-based reinforcement learning [61, 21, 62], which learns an approximate dynamics model of the environment. Given a downstream reward function, task-optimal behavior can be obtained by “planning” via model rollouts [50, 58, 39, 45]. Typically, model rollouts are generated autoregressively, conditioning each step on generation from the previous step. In practice, however, autoregressive generation suffers from compounding error [31, 1, 26], which arises when small, one-step approximation errors accumulate over long horizons. This leads to rollout trajectories that diverge from real trajectories, limiting many model-based RL methods to short-horizon, low-dimensional problems.
@@
An alternative class of algorithms based on successor features (SFs) has emerged as a potential approach to transferable decision-making [3, 2]. Successor features represent the discounted sum of features for a given policy. Assuming a linear correspondence between features and rewards, policy evaluation under new rewards reduces to a simple linear regression problem. Notably, by directly predicting long-term outcomes, SFs avoid autoregressive rollouts and hence compounding error. However, the notion of successor features is deeply tied to the choice of a particular policy. This policy dependence hinders the recovery of optimal policies for various downstream tasks. Current approaches to circumvent policy dependence either maintain a set of policies and select the best one during inference [3] or randomly sample reward vectors and make conditional policy improvements [5, 54, 55]. Nevertheless, a turnkey solution to transfer remains a desirable goal.
@@
In this work, we propose a new class of models—Distributional Successor Features for Zero-Shot Policy Optimization (DiSPOs)—that are rapidly transferable across reward functions while avoiding compounding errors. Rather than modeling the successor features under a particular policy, DiSPOs model the distribution of successor features under the behavior policy, effectively encoding all possible outcomes that appear in the dataset at each state. Crucially, by representing outcomes as successor features, we enjoy the benefit of zero-shot outcome evaluation after solving a linear reward regression problem, without the pitfall of compounding error. In addition to the outcome model, DiSPOs jointly learn a readout policy that generates an action to accomplish a particular outcome. Together, these models enable zero-shot policy optimization [55] for arbitrary rewards without further training: at test time, we simply perform a linear regression, select the best in-distribution outcome, and query the readout policy for an action to realize it. DiSPOs are a new class of world models as they essentially capture the dynamics of the world and can be used to plan for optimal actions under arbitrary rewards, without facing the pitfalls of compounding error.
@@
Since multiple outcomes can follow from a particular state, and multiple actions can be taken to achieve a particular outcome, both the outcome distribution and the policy require expressive model classes to represent. We provide a practical instantiation of DiSPOs using diffusion models [23, 47] and show that under this parameterization, policy optimization can be cast as a variant of guided diffusion sampling [13]. We validate the transferability of DiSPOs across a suite of long-horizion simulated robotics domains and further show that DiSPOs provably converge to “best-in-data” policies. With DiSPOs, we hope to introduce a new way for the research community to envision transfer in reinforcement learning, and to think of alternative ways to address the challenges of world modeling.
@@ 2 Related Work
Model-Based RL To enable transfer across rewards, model-based RL learns one-step (or multi-step) dynamics models via supervised learning and use them for planning  [11, 38, 39, 58] or policy optimization  [12, 50, 26, 64, 20]. These methods typically suffer from compounding error, where autoregressive model rollouts lead to large prediction errors over time  [1, 31]. Despite improvements to model architectures  [20, 25, 1, 32, 65] and learning objectives [27], modeling over long horizons without compounding error remains an open problem. DiSPOs instead directly model cumulative long-term outcomes in an environment, avoiding autoregressive generation while remaining transferable.
@@
Successor Features Successor features achieve generalization across rewards by modeling the accumulation of features (as opposed to rewards in model-free RL)  [3, 2]. With the assumption that rewards are linear in features, policy evaluation under new rewards reduces to a linear regression problem. A key limitation of successor features is their inherent policy dependence, as they are defined as the accumulated features when acting according to a particular policy. This makes extracting optimal policies for new tasks challenging.
@@
To circumvent this policy dependence, generalized policy improvement [3, 2] maintains a discrete set of policies and selects the highest valued one to execute at test time, limiting the space of available policies for new tasks. Universal SF [6] and Forward-Backward Representations [54, 55] randomly sample reward weights $z$ and jointly learn successor features and policies conditioned on $z$. The challenge lies in achieving coverage over the space of all possible policies through sampling of $z$, resulting in potential distribution shifts for new problems. RaMP [9] learns a successor feature predictor conditioned on an initial state and a sequence of actions. Transfer requires planning by sampling actions sequences, which becomes quickly intractable over horizon. In contrast, DiSPOs avoid conditioning on any explicit policy representation by modeling the distribution of all possible outcomes represented in a dataset, and then selecting actions corresponding to the most desirable long-term outcome.
@@
Distributional Successor Measure (DSM) [59] is a concurrent work that learns a distribution over successor representations using tools from distributional RL [4]. Importantly, DSM models the distributional successor measure of a particular policy, where the stochasticity stems from the policy and the dynamics. This makes it suitable for robust policy evaluation but not for transferring to arbitrary downstream tasks. In contrast, DiSPOs model the distribution of successor feature outcomes in the dataset (i.e., the behavior policy), where the distribution stems from the range of meaningfully distinct long-term outcomes. This type of modeling allows DiSPOs to extract optimal behavior for arbitrary downstream tasks, while DSMs suffer from the same policy dependence that standard successor feature-based methods do.
@@ 3 Preliminaries
Successor features [2] generalize the notion of a value function from task-specific rewards to task-agnostic features. Given a state feature function $\phi:S\rightarrow\mathbb{R}^{d},$ the successor feature of a policy is defined as $\psi^{\pi}(s)=\mathbb{E}_{\pi,\mathcal{T}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}\phi(s_{i})\mid s_{1}=s\right]$. Suppose rewards can be linearly expressed by the features, i.e. there exists $w\in\mathbb{R}^{n}$ such that $R(s)=w^{\top}\phi(s),$ then the value function for the particular reward can be linearly expressed by the successor feature $V^{\pi}(s)=w^{\top}\psi^{\pi}(s).$ Hence, given the successor feature $\psi^{\pi}$ of a policy $\pi$, we can immediately compute its value under any reward once the reward weights $w$ are known. Analogous to value functions, successor features also admit a recursive Bellman identity $\psi^{\pi}(s):=\phi(s)+\gamma\mathbb{E}_{\pi}\left[\psi^{\pi}(s^{\prime})\right],$ allowing them to be estimated using dynamic programming [3]. In this paper, we also refer to the discounted sum of features along a trajectory as a successor feature. In this sense, a successor feature represents an outcome that is feasible under the dynamics and can be achieved by some policy.
Diffusion Models DiSPOs rely on expressive generative models to represent the distribution of successor features. Diffusion models [23, 47] are a class of generative models where data generation is formulated as an iterative denoising process. Specifically, DDPM [23] consists of a forward process that iteratively adds Gaussian noise to the data, and a corresponding reverse process that iteratively denoises a unit Gaussian to generate samples from the data distribution. The reverse process leverages a neural network estimating the score function of each noised distribution, trained with a denoising score matching objective [49]. In addition, one can sample from the conditional distribution $p(x|y)$ by adding a guidance $\nabla_{x}\log p(y|x)$ to the score function in each sampling step [13]. As we show in Sec. 4.3, guided diffusion enables quick selection of optimal outcomes from DiSPOs.
@@
4 Distributional Successor Features for Zero-Shot Policy Optimization
We introduce the framework of DiSPOs as a scalable approach to the transfer problem described in Section 3, with the goal of learning from an unlabeled dataset to quickly adapt to any downstream task specified by a reward function. We start by relating the technical details behind learning DiSPOs in Section 4.1, followed by explaining how DiSPOs can be used for efficient multi-task transfer in Section 4.2. Finally, we describe a practical instantiation of DiSPOs in Section 4.3.
4.1 Learning Distributional Successor Features of the Behavior Policy
To transfer and obtain optimal policies across different reward functions, generalist decision-making agents must model the future in a way that permits the evaluation of new rewards and new policies. To this end, DiSPOs adopt a technique based on off-policy dynamic programming to directly model the distribution of cumulative future outcomes, without committing to a particular reward function $r(\cdot)$ or policy $\pi$. Fig. 2 illustrates the two components in DiSPOs, and we describe each below.
(1) Outcome model: for a particular a state feature function $\phi(s)$, DiSPOs model the distribution of successor features $p(\psi|s)$ over all paths that have coverage in the dataset. In deterministic MDPs, each successor feature $\psi$ (discounted sum of features $\psi=\sum_{t}\gamma^{t-1}\phi(s_{t})$) can be regarded as an “outcome". When the state features are chosen such that reward for the desired downstream task is a linear function of features, i.e., there exists $w\in\mathbb{R}^{n}$ such that $r(s)=w^{\top}\psi(s)$  [42, 43, 66, 56, 9, 3], the value of each outcome can be evaluated as $w^{\top}\psi$. That is, knowing $w$ effectively transforms the distribution of outcomes $p(\psi|s)$ into a distribution of task-specific values (sum of rewards) $p(R|s)$. Notably, since $w$ can be estimated by regressing rewards from features, distributional evaluation on a new reward function boils down to a simple linear regression.
As in off-policy RL, the outcome distribution $p(\psi|s)$ in DiSPOs can be learned via an approximate dynamic programming update, which is similar to a distributional Bellman update [4]:
@@ 4.1 Learning Distributional Successor Features of the Behavior Policy
(2) Readout policy: Modeling the distribution of future outcomes in an environment is useful only when it can be realized in terms of actions that accomplish particular outcomes. To do so, DiSPOs pair the outcome model with a readout policy $\pi(a|s,\psi)$ that actualizes a desired long-term outcome $\psi$ into the action $a$ to be taken at state $s$. Along with the outcome model $p_{\theta}(\psi|s)$, the readout policy $\pi_{\rho}(a|s,\psi)$ can be optimized via maximum-likelihood estimation:
@@
Figure 2: DiSPOs for a simple environment. Given a state feature function $\phi$, DiSPOs learn a distribution of all possible long-term outcomes (successor features $\psi$) in the dataset $p(\psi|s)$, along with a readout policy $\pi(a|s,\psi)$ that takes an action $a$ to realise $\psi$ starting at state $s$.
@@
4.2 Zero-Shot Policy Optimization with Distributional Successor Features
To synthesize optimal policies for novel downstream reward functions using DiSPOs, two sub-problems must be solved: (1) inferring the suitable linear reward weights $w_{r}$ for a particular reward function from a set of $(s,r)$ tuples and (2) using the inferred $w_{r}$ to select an optimal action $a^{*}$ at a state $s$. We discuss each below.
Inferring task-specific weights with linear regression. As noted, for any reward function $r(s)$, once the linear reward weights $w_{r}$ are known (i.e., $r(s)=w_{r}^{T}\phi(s)$), the distribution of returns in the dataset $p(R|s)$ is known through linearity. However, in most cases, rewards are not provided in functional form, making $w_{r}$ unknown a priori. Instead, given a dataset of $\mathcal{D}=\{(s,r)\}$ tuples, $w_{r}$ can be obtained by solving a simple linear regression problem $\arg\min_{w_{r}}\frac{1}{|\mathcal{D}|}\sum_{(s,r)\in\mathcal{D}}\|w_{r}^{T}\psi(s)-r\|_{2}^{2}$.
Generating task-specific policies via distributional evaluation. Given the inferred $w_{r}$ and the corresponding future return distribution $p(R|s)$ obtained through linear scaling of $p(\psi|s)$, the optimal action can be obtained by finding the $\psi$ with the highest possible future return that has sufficient data-support: